# Approaching Stable Quark Matter

### v1

@@ 4 Bag parameter from Lattice QCD calculation of isospin-dense matter
$Z_{\rm I}(\mu)$ $=\int\mathcal{D}[A]\det\mathcal{D}(-\mu)\det\mathcal{D}(\mu)e^{-S_{G}}=\int\mathcal{D}[A]\,\left|\det\mathcal{D}(\mu)\right|^{2}e^{-S_{G}}\leavevmode\nobreak\ ,$ (4.1)
where $A$ is the gluon field, $S_{G}$ is the gluon field action, and $\mathcal{D}(\mu)=\not{D}+m_{q}-\mu\gamma_{0}$ is the Dirac operator. It is clear that $Z_{\rm I}$ is free of the sign problem as it only depends on the norm $|\det\mathcal{D}|$, thus allowing one to perform LQCD simulations for a wide range of $\mu$ values.
Recently, the NPLQCD collaboration performed a simulation on a large isospin-dense system up to $\mu\sim 1.6$ GeV and made their data publicly available in Ref. [18] (their earlier simulations can be found in Ref. [17]). Notably, the isospin-dense matter system shares the same bag parameter $B$ as the baryon-dense matter system. We will analyze the LQCD data for the isospin-dense system to obtain the value of or the bound on $B$.
@@
As first pointed out in Ref. [22], comparing the LQCD data to the pQCD calculations at finite density will shed much light on our understanding of QCD. While Ref. [22] focuses on the role played by the CS condensates, we are interested in the inference of the bag parameter through comparing the LQCD and pQCD calculations, which would in turn imply the stability of the quark matter. Such is possible since $B$ is universal to both the baryon and isospin-dense matter, an important point that allows one to study the latter free of the sign problem with LQCD to infer about the former, as we have mentioned earlier. The pQCD model to fit the isospin-dense LQCD data is given by
$p(\mu,X,B)=p_{\rm pQCD}(\mu,X)\,+\,p_{\rm CS}(\mu,X)\,-\,B\leavevmode\nobreak\ ,$ (4.2)
@@
$\epsilon=-p+\mu\frac{\partial p}{\partial\mu},\leavevmode\nobreak\ \qquad c_{s}^{2}=\frac{\partial p/\partial\mu}{\partial\epsilon/\partial\mu}\leavevmode\nobreak\ ,$ (4.3)
where the details of $p_{\rm pQCD}$ and $p_{\rm CS}$ can be found in Appendix A and Appendix C, respectively  [^3] We note that in this work $p_{\rm QCD}$ is calculated up to the three-loop level, while $p_{\rm CS}$ is calculated up to the two-loop level diagrammatically. One can make the predictions more consistent by improving the calculations of $p_{\rm CS}$ to a higher-loop order in the future.. There are two independent model parameters to be fit: $X$ and $B$.
@@
Superficially, the overall degrees of freedom of the data seem to be given by $300$($\mu$ bins)×$\times 3$ (observables) $=900$. However, as pointed out in Ref. [18], the correlation among the neighboring $\mu$ bins and among the different observables greatly reduces the effective degrees of freedom of the data. This is reflected in the large hierarchy among the eigenvalues of the covariance matrix $\Sigma$ (which we show along with the associated eigen-observables in Figure 9 of Appendix D), and thus we should truncate the fit at a certain number of the effective degrees of freedom $k$ to avoid overfitting. We note that the total $\chi^{2}$ will be dominated by the last few degrees of freedom, which carry the smallest eigenvalues of $\Sigma$. These degrees of freedom, assuming we are not overfitting the data, will provide the most stringent constraints on the parameter space.
For each $\mu_{\rm start}$, we determine the values of $k$ by requiring the best-fit point to sit just within the $1\sigma$ and $90\%$ confidence level (CL) contours for $k-2$ degrees of freedom. For $\mu_{\rm start}=1000$ MeV, one has $k=49$ and $53$ for the $1\sigma$ and $90\%$ CL best-fit conditions, respectively, whose benchmark parameters are $(X,B^{1/4})=(1.66,0\leavevmode\nobreak\ {\rm MeV})$ and $(X,B^{1/4})=(1.78,0\leavevmode\nobreak\ {\rm MeV})$. The LQCD data prefers a zero bag parameter based on the prior $B\geq 0$. This is mainly driven by the smaller energy density $\epsilon(\mu)$ of the LQCD data compared to the pQCD predictions (see the right panel of Figure 5). If we require a smaller confidence level that hints at overfitting the data, one has a smaller number of effective degrees of freedom $k$.
After determining the best-fit points and the corresponding values of $k$, we trace out the $90\%$ CL contour on the $(X,B^{1/4})$ plane, which will provide the bounds on both $X$ and $B$. In Figure 5, we show the $c_{s}^{2}$, $p/p_{\rm id}$, and $\epsilon/\epsilon_{\rm id}$ distributions of the two best-fit points (red and green solid lines) as well as those of the benchmark $(X,B^{1/4})=(1.66,161\leavevmode\nobreak\ {\rm MeV})$ (red dashed line), which sits on the $90\%$ CL contour around the $k=49$ best-fit point. For illustration, we further show the predictions of two other benchmarks: $(X,B^{1/4})=(1.66,200\leavevmode\nobreak\ {\rm MeV})$ (red dotted line) and $(X,B^{1/4})=(1.66,250\leavevmode\nobreak\ {\rm MeV})$ (red dot-dashed line), with the latter clearly showing a discrepancy from the LQCD data. Note that $c_{s}^{2}$ does not rely on $B$ and thus the red lines completely overlap with one another in the left panel.
Figure 5: The $c_{s}^{2}$, $p/p_{\rm id}$, and $\epsilon/\epsilon_{\rm id}$ distributions of the best-fit points (red and green solid lines) with $\mu_{\rm start}=1000$ MeV and $k=49,53$, whose benchmark parameters are $(X,B^{1/4})=(1.66,0\leavevmode\nobreak\ {\rm MeV})$ and $(1.78,0\leavevmode\nobreak\ {\rm MeV})$. Also shown are the distributions of the benchmarks $(X,B^{1/4})=(1.66,161\leavevmode\nobreak\ {\rm MeV})$ (red dashed line), which sits on the $90\%$ CL contour around the $1\,\sigma$ ($k=49$) best-fit point, $(X,B^{1/4})=(1.66,200\leavevmode\nobreak\ {\rm MeV})$ (red dotted line), and $(1.66,250\leavevmode\nobreak\ {\rm MeV})$ (red dot-dashed line).
@@ 5 Inferred quark matter stability
Along the bag parameter $B$ direction, the current LQCD data prefers a zero value of $B$, which can be traced back to the $\epsilon(\mu)$ distribution shown in the right panel of Figure 5 that is lower than the naively anticipated value of $B^{1/4}=\mathcal{O}(100\,\mbox{MeV})$ or the QCD scale. Interestingly, the 90% CL constraints on the bag parameter from the $\mbox{LQCD}_{\rm I}$ data exert the limit $B^{1/4}\lesssim 160$ MeV, which is close to but above the threshold (80-120 MeV for $X\in[1.5,2.1]$) for stable quark matter. Therefore, though there is no conclusive statement on the stability of quark matter, one important message from Figure 6 is that the combined information from pQCD and $\mbox{LQCD}_{\rm I}$ is approaching the answer to the question of quark matter stability.
@@
    If the benchmark falls within the blue shaded region in Figure 6, then the existence of any type of stable quark matter is excluded by the pQCD calculation.
@@
    If the benchmark falls within the remaining white region in Figure 6, then it implies that either the 2+1- or both types of quark matter are more stable than ordinary nuclei based on the pQCD calculation.
Figure 6: A superposition of the allowed $X$-$B^{1/4}$ parameter space for stable quark matter from the right panel of Figure 2 based on pQCD calculations and the constraints from $\mbox{LQCD}_{\rm I}$ simulation data. The red and green triangle points are the best-fit points for $k=49$ (the total $\chi^{2}$ at 1 $\sigma$) and $k=53$ (the total $\chi^{2}$ at 90% CL). The red and green solid lines are the 90% CL contours around the best-fit points. The red and green dot-dashed lines are the “projected limits” with the same values of $k=49$ (red) and $k=53$ (green) with a decrease in the LQCD uncertainties by a factor of $N_{\rm scale}=\sqrt{10}$. $\mu_{\rm start}=1000$ MeV is used for the LQCD data analysis.
@@ 6 General bounds on quark matter stability from thermodynamics
In this section, we adopt a complementary approach to the main LQCD+pQCD framework of this work to explore the quark matter stability by applying the thermodynamic bounds on the quark matter equation of state (EOS) following the methodology outlined in Ref. [23]. In this approach, one imposes a few basic thermodynamic requirements that can also provide a bound on the relation between the quark matter energy per baryon $\epsilon/n_{\text{B}}$ and the bag parameter $B$, and hence the stability of quark matter. The requirements are: 1. the pressure $p(\mu_{\text{B}})$ being a continuous function of $\mu_{\text{B}}$; 2. the thermodynamic stability or concavity of the grand potential; 3. the causality bound that requires the speed of sound be smaller than the speed of light; and 4. the low-density reference point with $(\mu_{L},p_{L})$ and high-density reference point with $(\mu_{H},p_{H})$ ($\mu_{L}$ and $\mu_{H}$ are baryon chemical potentials) being connected by an integral relation. For the details of the framework, we refer to Appendix E and Appendix F.
@@
Figure 7: General thermodynamic bounds on the $B^{1/4}$$(\epsilon/n_{\text{B}})$ relation for different high-density reference benchmarks, based on the 2+1-flavor pQCD calculations. We also show the current (vertical red solid line) and projected (vertical red dot-dashed line) LQCD sensitivities for $X=1.6$, $\mu_{\rm start}=1000$ MeV, and $k=49$ (see Figure 6). The horizontal black dashed line represents $\epsilon/n_{\text{B}}=930$ MeV.
In Figure 7, we show such constraints on the $B^{1/4}$$(\epsilon/n_{\text{B}})$ parameter space dervied from the general thermodynamic requirements using three example 2+1-flavor benchmarks with $(\mu_{H},X)=(2.0\leavevmode\nobreak\ {\rm GeV},1.6)$, $(1.5\leavevmode\nobreak\ {\rm GeV},1.6)$, and $(1.0\leavevmode\nobreak\ {\rm GeV},1.6)$. One can see that the constraints are more stringent for a lower value of $\mu_{H}$. On the other hand, a lower value of $\mu_{H}$ implies a larger uncertainty in $p_{H}$ based on the pQCD calculation, so the derived constraints are less reliable. For the three benchmarks, the QCD gauge couplings are $\alpha_{s}(2X\mu_{H}/3)\approx 0.29,0.34,0.49$, respectively, and thus one could trust the constraints derived from the first two benchmarks (green and orange shaded regions) but not (at least not completely) the last one. For illustration, we also show the constraining contour in terms of a blue dot-dashed line for the third benchmark. The intersections between the constraining contours and the black dashed line ($\epsilon/n_{\text{B}}=930\leavevmode\nobreak\ \mbox{MeV}$) are $B^{1/4}=165$ MeV (215 MeV) for the green (orange) curves. This means that for a bag parameter above these values, the existence of stable 2+1-flavor quark matter is inconsistent with the general requirements of thermodynamic equilibrium derived with the two corresponding high-density reference points. On the other hand, both stable and unstable quark matter could exist in the white allowed region, so no conclusive answers can be obtained from this more general approach.
For comparison, we also show the current and projected LQCD sensitivities on $B$ with $X=1.6$, $\mu_{\rm start}=1000$ MeV, and $k=49$ on this plot (see Figure 6). Both approaches provide comparable constraints on $B$, although the LQCD+pQCD framework can be improved with future LQCD studies.

### v3

@@ 4 Bag parameter from Lattice QCD calculation of isospin-dense matter
$Z_{\rm I}(\mu)$ $=\int\mathcal{D}[A]\det\mathcal{D}(-\mu)\det\mathcal{D}(\mu)e^{-S_{G}}=\int\mathcal{D}[A]\,\left|\det\mathcal{D}(\mu)\right|^{2}e^{-S_{G}}\penalty 10000\ ,$ (4.1)
[…]
$p_{\rm B}(\mu_{\rm B})\leq p_{\rm I}\left(\mu_{\rm I}=\frac{2\mu_{\rm B}}{N_{c}}\right)\penalty 10000\ .$ (4.2)
This has been investigated in Ref. [39] using the LQCD dataset from Ref. [27] (we will discuss it more later). Following their methodology but using the more updated LQCD data from Ref. [28], we find that the resulting bound $p_{\rm B}(\mu_{\rm B}=930\penalty 10000\ {\rm MeV})\leq(210\penalty 10000\ {\rm MeV})^{4}$], though robust, is much looser than the pressure values derived using pQCD for the $X$ values we consider $p_{\rm B}(\mu_{\rm B}=930\penalty 10000\ {\rm MeV})\leq(150\penalty 10000\ {\rm MeV})^{4}$], and thus we will not discuss them further for our purpose of studying quark matter stability. Nevertheless, this bound could be useful for the investigation of quark matter scenarios beyond the capability of pQCD.
Recently, the NPLQCD collaboration performed a simulation on a large isospin-dense system up to $\mu\sim 1.6$ GeV at $T\sim\mathcal{O}(10)\penalty 10000\ {\rm MeV}$ (which should be low enough to be considered as “cold”) and made their data publicly available in Ref. [28] (their earlier simulations can be found in Ref. [27]). Notably, the isospin-dense matter system shares the same bag parameter $B$ as the baryon-dense matter system. We will analyze the LQCD data for the isospin-dense system to obtain the value of or the bound on $B$.
@@
As first pointed out in Ref. [32], comparing the LQCD data to the pQCD calculations at finite density will shed much light on our understanding of QCD. While Ref. [32] focuses on the role played by the CS condensates, we are interested in the inference of the bag parameter through comparing the LQCD and pQCD calculations, which would in turn imply the stability of the quark matter. Such is possible since $B$ is universal to both the baryon and isospin-dense matter, an important point that allows one to study the latter free of the sign problem with LQCD to infer about the former, as we have mentioned earlier. Given the crossover nature of the phase transition in the isospin-dense system, one might question the validity of introducing $B$ and modeling the system using pQCD; however, as we will discuss later, since we only analyze the data with $\mu>1000$ GeV (an alternative analysis of data with $\mu>800$ MeV is given in Appendix D), perturbativity should be valid while the vacuum description with $B$ is asymptotically correct. The pQCD model to fit the isospin-dense LQCD data is given by
$p(\mu,X,B)=p_{\rm pQCD}(\mu,X)\,+\,p_{\rm CS}(\mu,X)\,-\,B\penalty 10000\ ,$ (4.3)
@@
$\epsilon=-p+\mu\frac{\partial p}{\partial\mu},\penalty 10000\ \qquad c_{s}^{2}=\frac{\partial p/\partial\mu}{\partial\epsilon/\partial\mu}\penalty 10000\ ,$ (4.4)
where the details of $p_{\rm pQCD}$ and $p_{\rm CS}$ can be found in Appendix A and Appendix C, respectively. [^5] We note that in this work $p_{\rm QCD}$ is calculated up to the three-loop level, while $p_{\rm CS}$ is calculated up to the two-loop level diagrammatically. One can make the predictions more consistent by improving the calculations of $p_{\rm CS}$ to a higher-loop order in the future. There are two independent model parameters to be fit: $X$ and $B$.
@@
[…]
For each $\mu_{\rm start}$, we determine the values of $k$ by requiring the best-fit point to sit just within the $1\sigma$ and $90\%$ confidence level (CL) contours for k2k-2 degrees of freedom. For $\mu_{\rm start}=1000$ MeV, one has $k=49$ and $53$ for the $1\sigma$ and $90\%$ CL best-fit conditions, respectively, whose benchmark parameters are $(X,B^{1/4})=(1.66,0\penalty 10000\ {\rm MeV})$ and $(X,B^{1/4})=(1.78,0\penalty 10000\ {\rm MeV})$. The LQCD data prefers a zero bag parameter based on the prior $B\geq 0$. This is mainly driven by the smaller energy density $\epsilon(\mu)$ of the LQCD data compared to the pQCD predictions (see the right panel of Figure 5). If we require a smaller confidence level that hints at overfitting the data, one has a smaller number of effective degrees of freedom $k$.
After determining the best-fit points and the corresponding values of $k$, we trace out the $90\%$ CL contour on the $(X,B^{1/4})$ plane, which will provide the bounds on both $X$ and $B$. In Figure 5, we show the $c_{s}^{2}$, $p/p_{\rm id}$, and $\epsilon/\epsilon_{\rm id}$ distributions of the two best-fit points (red and green solid lines) as well as those of the benchmark $(X,B^{1/4})=(1.66,161\penalty 10000\ {\rm MeV})$ (red dashed line), which sits on the $90\%$ CL contour around the $k=49$ best-fit point. For illustration, we further show the predictions of two other benchmarks: $(X,B^{1/4})=(1.66,200\penalty 10000\ {\rm MeV})$ (red dotted line) and $(X,B^{1/4})=(1.66,250\penalty 10000\ {\rm MeV})$ (red dot-dashed line), with the latter clearly showing a discrepancy from the LQCD data. Note that $c_{s}^{2}$ does not rely on $B$ and thus the red lines completely overlap with one another in the left panel.
Figure 5: The $c_{s}^{2}$, $p/p_{\rm id}$, and $\epsilon/\epsilon_{\rm id}$ distributions of the best-fit points (red and green solid lines) with $\mu_{\rm start}=1000$ MeV and $k=49,53$, whose benchmark parameters are $(X,B^{1/4})=(1.66,0\penalty 10000\ {\rm MeV})$ and $(1.78,0\penalty 10000\ {\rm MeV})$. Also shown are the distributions of the benchmarks $(X,B^{1/4})=(1.66,161\penalty 10000\ {\rm MeV})$ (red dashed line), which sits on the $90\%$ CL contour around the $1\,\sigma$ ($k=49$) best-fit point, $(X,B^{1/4})=(1.66,200\penalty 10000\ {\rm MeV})$ (red dotted line), and $(1.66,250\penalty 10000\ {\rm MeV})$ (red dot-dashed line).
@@ 5 Inferred quark matter stability
Along the bag parameter $B$ direction, the current LQCD data prefers a zero value of $B$, which can be traced back to the $\epsilon(\mu)$ distribution shown in the right panel of Figure 5 that is lower than the naively anticipated value of $B^{1/4}=\mathcal{O}(100\,\mbox{MeV})$ or the QCD scale. Interestingly, the 90% CL constraints on the bag parameter from the $\mbox{LQCD}_{\rm I}$ data exert the limit $B^{1/4}\lesssim 160$ MeV, which is close to but above the threshold (80-120 MeV for $X\in[1.5,2.1]$) for stable quark matter. Therefore, though there is no conclusive statement on the stability of quark matter (at least for the 2-flavor one), one important message from Figure 6 is that the combined information from pQCD and $\mbox{LQCD}_{\rm I}$ is approaching the answer to the question of quark matter stability.
In Figure 6, we also show existing lowered bounds on the bag parameter from the Gell-Mann, Oakes, Renner (GMOR) relation [18] and the QCD sum rules, which we briefly mentioned in Section 1. Note that the term $\sum_{q}m_{q}\bar{q}q$ in the trace anomaly defined in Eq. (1.1) is independent of the renormalization scale; the same is true for the sum of the two remaining ($\beta(g)$ and $\gamma_{m}(g)$) terms [42].
For the contribution of $B_{m}\equiv\langle\sum_{q}m_{q}\bar{q}q\rangle$ to $B$, we have a reliable answer from the GMOR relation [18], especially for the two light ($u$ and $d$) quarks
$\frac{f_{\pi}^{2}m_{\pi}^{2}}{4}=-m_{\ell}\langle\overline{\ell}\ell\rangle\penalty 10000\ ,$ (5.1)
where $f_{\pi}=130$ MeV and $m_{\ell}=(m_{u}+m_{d})/2$, which in the $\overline{\rm MS}$ scheme at 2 GeV using the $SU(2)$ chiral perturbation theory gives $\langle\overline{\ell}\ell\rangle^{\overline{\rm MS}}(2\penalty 10000\ {\rm GeV})=-(272\pm 2\penalty 10000\ {\rm MeV})^{3}$ [21] and is consistent with the lattice results [22]. On the other hand, since $m_{s}$ is relatively heavy compared to $m_{u}$ and $m_{d}$, there is a greater disagreement among different measurements. For instance, Ref. [24] has summarized a couple of different measurements, from which we especially quote
[…]
On the other hand, there is a large uncertainty in the second contribution of $B_{G}\equiv\langle\beta\,GG/2g+\sum_{q}m_{q}\,\gamma_{m}\,\bar{q}q\rangle$ to $B$. See for example Ref. [23] for an overview of various different measurements of the gluon condensate $\langle\alpha_{s}GG\rangle$. Some values of $\langle\alpha_{s}GG\rangle$ (not renormalization-scale-independent) obtained from the sum-rule approach are in conflict with the upper bound on $B$ extracted from the LQCDI data in Section 4. As a result, we remain agnostic to this second contribution to $B$ and only assume $B_{G}\geq 0$. Accordingly, we have a trustable lower bound on $B$
$B\geq B_{{\rm min},2+1}\equiv-\frac{1}{4}\sum_{q=u,d,s}\langle m_{q}\overline{q}q\rangle^{\overline{\rm MS}}(2\penalty 10000\ {\rm GeV})=\begin{cases}(153\begin{subarray}{c}+6.0\\ -6.0\end{subarray}\penalty 10000\ {\rm MeV})^{4}&\penalty 10000\ [\text{Lattice}]\\[4.30554pt] (136\pm 5.0\penalty 10000\ {\rm MeV})^{4}&\penalty 10000\ [\text{Sum Rule Average}]\end{cases}\penalty 10000\ .$ (5.3)
For the 2-flavor system, since the $\langle m_{s}\overline{s}s\rangle$ condensate may not be restored [44], one could only consider the contributions from the $u$- and $d$-quark chiral condensates to the bag parameter $B$. In this case, one would have a lower value for the lower bound
$B\geq B_{{\rm min},2}\equiv-\frac{1}{4}\sum_{q=u,d}\langle m_{q}\overline{q}q\rangle^{\overline{\rm MS}}(2\penalty 10000\ {\rm GeV})=(76.8\begin{subarray}{c}+0.9\\ -2.0\end{subarray}\penalty 10000\ {\rm MeV})^{4}\penalty 10000\ \penalty 10000\ [\text{GMOR}]\penalty 10000\ .$ (5.4)
As can be seen from the left panel of Figure 6, the lower bound on $B$ from Eq. (5.3) has entirely excluded the existence of the stable 2+1-flavor quark matter in the $X$ range of our interest. Given that in order for the 2+1-flavor quark matter to be stable, the trend of the stability constraint contour (the blue solid line) requires a smaller value of $B$ for a smaller value of $X$, one can draw the interesting conclusion that there is no stable 2+1-flavor quark matter. We also note that the $\langle m_{s}\bar{s}s\rangle$ condensate may not be restored in the isospin-dense matter. In this case, the upper bound on $B$ from LQCDI can be treated as a more conservative bound for the 2+1-flavor quark matter.
In the right panel, we compare the 2-flavor quark matter stability constraint contour (the orange solid line) with the LQCDI constraints and the lower bound $B_{{\rm min},2}$ from the GMOR relation. One can see that there is only a narrow region of $X$ and $B$ that allows stable 2-flavor quark matter. It will be interesting to see if the future determination of $B$ falls into this allowed region.
Figure 6: Left panel: A superposition of the allowed $X$-$B^{1/4}$ parameter space for the 2+1-flavor stable quark matter from the right panel of Figure 2 based on pQCD calculations and the constraints from $\mbox{LQCD}_{\rm I}$ simulation data. The red and green triangle points are the best-fit points for $k=49$ (the total $\chi^{2}$ at 1 $\sigma$) and $k=53$ (the total $\chi^{2}$ at 90% CL). The red and green solid lines are the 90% CL contours around the best-fit points. The red and green dot-dashed lines are the “projected limits” with the same values of $k=49$ (red) and $k=53$ (green) with a decrease in the LQCD uncertainties by a factor of $N_{\rm scale}=\sqrt{10}$. $\mu_{\rm start}=1000$ MeV is used for the LQCD data analysis. The lower bounds from Eq. (5.3) are also shown in the plot. Right panel: Same as the left panel but for the 2-flavor stable quark matter, with the lower bound from Eq. (5.4) also shown in the plot.
@@
    If the benchmark falls within the blue shaded region in the left panel or the orange shaded region in the right panel of Figure 6, then the existence of any type of stable quark matter is excluded by the pQCD calculation.
@@
    If the benchmark falls within the remaining white region in the right panel of Figure 6, then it implies that the 2-flavor quark matter is more stable than ordinary nuclei based on the pQCD calculation.
@@ 6 General bounds on quark matter stability from thermodynamics
In this section, we adopt a complementary approach to the main LQCD+pQCD framework of this work to explore the quark matter stability by applying the thermodynamic bounds on the quark matter equation of state (EOS) following the methodology outlined in Ref. [33] (also considered in Ref. [39]). In this approach, one imposes a few basic thermodynamic requirements that can also provide a bound on the relation between the quark matter energy per baryon $\epsilon/n_{\text{B}}$ and the bag parameter $B$, and hence the stability of quark matter. The requirements are: 1. the pressure $p(\mu_{\text{B}})$ being a continuous function of $\mu_{\text{B}}$; 2. the thermodynamic stability or concavity of the grand potential; 3. the causality bound that requires the speed of sound be smaller than the speed of light; and 4. the low-density reference point with $(\mu_{L},p_{L})$ and high-density reference point with $(\mu_{H},p_{H})$ ($\mu_{L}$ and $\mu_{H}$ are baryon chemical potentials) being connected by an integral relation. For the details of the framework, we refer to Appendix E and Appendix F.
@@
Figure 7: General thermodynamic bounds on the $B^{1/4}$$(\epsilon/n_{\text{B}})$ relation for different high-density reference benchmarks, based on the 2-flavor pQCD calculations. We also show the current (vertical red solid line) and projected (vertical red dot-dashed line) LQCD sensitivities for $X=1.6$, $\mu_{\rm start}=1000$ MeV, and $k=49$ (see Figure 6). The horizontal black dashed line represents $\epsilon/n_{\text{B}}=930$ MeV. The lower (GMOR) bound from Eq. (5.4) is also shown in the plot.
In Figure 7, we show such constraints on the $B^{1/4}$$(\epsilon/n_{\text{B}})$ parameter space derived from the general thermodynamic requirements using three example 2-flavor benchmarks with $(\mu_{H},X)=(1.4\penalty 10000\ {\rm GeV},1.6)$, $(1.2\penalty 10000\ {\rm GeV},1.6)$, and $(1.0\penalty 10000\ {\rm GeV},1.6)$. One can see that the constraints are more stringent for a lower value of $\mu_{H}$. On the other hand, a lower value of $\mu_{H}$ implies a larger uncertainty in $p_{H}$ based on the pQCD calculation, so the derived constraints are less reliable. For the three benchmarks, the QCD gauge couplings are $\alpha_{s}(2X\mu_{H}/3)\approx 0.36,0.41,0.49$, respectively, and thus one could trust the constraints derived from the first two benchmarks (green and orange shaded regions) but not (at least not completely) the last one. For illustration, we also show the constraining contour in terms of a blue dot-dashed line for the third benchmark. The intersections between the constraining contours and the black dashed line ($\epsilon/n_{\text{B}}=930\penalty 10000\ \mbox{MeV}$) are $B^{1/4}=120$ MeV (140 MeV) for the green (orange) curves. This means that for a bag parameter above these values, the existence of stable 2-flavor quark matter is inconsistent with the general requirements of thermodynamic equilibrium derived with the two corresponding high-density reference points. On the other hand, both stable and unstable quark matter could exist in the white allowed region, so no conclusive answers can be obtained from this more general approach.
For comparison, we also show the current and projected LQCD sensitivities on $B$ with $X=1.6$, $\mu_{\rm start}=1000$ MeV, and $k=49$ on this plot (see Figure 6), as well as the lower bound from Eq. (5.4). Both approaches provide comparable constraints on $B$, although the LQCD+pQCD framework can be improved with future LQCD studies.